# Monster shocks, gamma-ray bursts and black hole quasi-normal modes from neutron-star collapse

### v1

@@ Abstract
Here $C_{\mu\nu\kappa\lambda}$ is the Weyl curvature tensor and $F^{\mu\nu}$ is the electromagnetic tensor; vectors $\bm{m}$ and $\bm{n}$ are conveniently chosen as $\bm{m}=(\hat{\bm{\theta}}+i\hat{\bm{\phi}})/\sqrt{2}$ and $\bm{n}=(\hat{\bm{t}}-\hat{\bm{r}})/\sqrt{2}$ (so that $\bm{m}$, $\bm{n}$, and $\bm{l}=(\hat{\bm{t}}+\hat{\bm{r}})/\sqrt{2}$ form an orthonormal null tetrad). Then, $\phi_{2}\propto E_{{\theta}}+iE_{\phi}$ represents two polarization states of the outgoing electromagnetic waves (in MHD, $E_{{\theta}}$ and $E_{{\phi}}$ correspond to the fast magnetosonic and Alfvén waves, respectively). We have verified that the dominant (quadrupole) component of $\psi_{4}$ observed in our simulation is consistent with the quasi-normal mode computed using qnm code [53]. We have also calculated the dominant (dipole) component of $\phi_{2}$, which also approximately matches the corresponding quasi-normal mode frequency (Fig. 4 shows the evolution of dominant spherical harmonics in ${\rm Im}\,\phi_{2}$ and $\psi_{4}$). Note however that $\phi_{2}$ was designed for vacuum electromagnetic fields, and so the oscillation of $\phi_{2}$ may not accurately represent the modulation of MHD outflow. The frequency of compressive modulations observed in Fig. 3 may be directly estimated as $\nu_{\rm mod}=v/\lambda$, where $v\lesssim c$ is the outflow speed and $\lambda\sim 7\,$km is the spatial modulation period.
@@
Gamma-ray burst. Our simulation demonstrates that the magnetospheric destruction during the neutron-star collapse involves strong dissipation and creates a powerful magnetized outflow with a characteristic duration $\sim 1$ ms. The hot outflow is launched behind the leading monster shock and has a quasi-periodic tail. The modulated tail is generated by the nascent black hole, as it rings down and quickly sheds most of its magnetic flux initially inherited from the neutron star. The modulation frequency lies naturally in the kilohertz band, suggesting a connection with recently reported kilohertz quasi-periodic oscillations in some GRBs [54].
@@
Note that the magnetosphere around the neutron star prior to collapse has a minute plasma density. Therefore, the explosion triggered by collapse is practically clean from baryons. The created $e^{\pm}$ plasma in the hot outflow is initially opaque to scattering. Most of the heat density $U$ is contained in trapped black body radiation $U\approx aT^{4}$ ($a$ is the radiation constant), as the photons far outnumber the $e^{\pm}$ pairs. The outflow expansion to large radii is not followed by our simulation, however its basic features can be predicted in analogy with the well-known “fireball” model for cosmological GRBs [55, 56]. The outflow Lorentz factor $\gamma$ will grow and its temperature $T$ will drop due to adiabatic cooling. Eventually, most of $e^{\pm}$ pairs annihilate, and the trapped photons are released, producing a GRB. The burst duration is set by the outflow duration, which lasts only a few ms after the collapse.
@@
The burst spectrum will peak at photon energies $\sim 3k_{B}T\gamma$, where $k_{B}$ is the Boltzmann constant. Simplest GRB models assume adiabatic outflows with no magnetic fields, and then $\gamma T\approx const=T_{0}$ [55]. In that case, the observed temperature is weakly changed during the outflow expansion and remains close to the initial temperature. Outflows predicted by our simulation are magnetically dominated, and their temperature may be affected by additional dissipation of magnetic energy at large radii. Regulation of the GRB spectral peak in dissipative outflows is discussed in [57].
@@
The energy of the GRB outflow is set by the pre-collapse energy of the neutron-star magnetosphere ${\cal E}_{m}\sim 10^{47}\mu_{33}^{2}\,$erg. During the collapse the magnetospheric energy is amplified by compression, and then most of it is ejected in the outflow. Assuming that $\sim 0.1$ of this energy is eventually emitted in the GRB, we can roughly estimate the GRB luminosity as $L\sim 0.1{\cal E}_{m}/1\,{\rm ms}\sim 10^{49}\,B_{15}^{2}\,$erg/s, where $B_{15}=B/10^{15}\,\rm G$ is the magnetic field near the neutron star prior to collapse. Note that the outflow (and hence the GRB) is anisotropic, but not strongly collimated.
@@
During its lifetime prior to collapse, the rotating magnetar produced a magnetic wind outside the dipole magnetosphere. The collapse launches a relativistic shock, which will continue to expand into the cold wind. This leading shock is expected to emit a fast radio burst (FRB) at radii $r\sim 10^{13}$ cm [8]. This suggests a mechanism for a delayed FRB from neutron star mergers, emitted together with the delayed GRB. However, the FRB can hardly escape through the surrounding shell of mass $M_{\rm ej}\sim 10^{-2}M_{\odot}$ ejected earlier by the merger [58, 59]. A similar problem is faced by the recently proposed FRB-GRB connection [60].
The massive shell is also a treat to the GRB predicted by our simulation. The GRB will not be blocked by $M_{\rm ej}$ if the magnetar collapse occurs with a sufficient delay $\Delta t\gtrsim 10$ hr after the merger, so that the massive ejecta expands to radius $R\sim 3\times 10^{14}\,(\Delta t/10{\rm\,hr})$ cm and its optical depth to gamma-rays $\tau\sim\kappa M_{\rm ej}/4\pi R^{2}$ drops below unity ($\kappa$ is the ejecta opacity to Compton scattering).
Acknowledgements.
The authors gratefully acknowledge insightful discussions with A. Philippov, V. Ravi, R. Sari, L. Sironi, J. Stone, S. Teukolsky and C. Thompson. ERM acknowledges support by the National Science Foundation under grants No. PHY-2309210 and AST-2307394. A.M.B. is supported by NSF AST 2009453, NASA 21-ATP21-0056 and Simons Foundation award No. 446228. B.R. is supported by the Natural Sciences & Engineering Research Council of Canada (NSERC) and by a grant from the Simons Foundation (MP-SCMPS-00001470). Simulations were performed on the NSF Frontera supercomputer under grant AST21006, and on Delta at the National Center for Supercomputing Applications (NCSA) through allocation PHY210074 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.

### v2

@@ 3.3 Black hole ring-down
Here $C_{\mu\nu\kappa\lambda}$ is the Weyl curvature tensor and $F^{\mu\nu}$ is the electromagnetic tensor; vectors $\bm{m}$ and $\bm{n}$ are conveniently chosen as $\bm{m}=(\hat{\bm{\theta}}+i\hat{\bm{\phi}})/\sqrt{2}$ and $\bm{n}=(\hat{\bm{t}}-\hat{\bm{r}})/\sqrt{2}$ (so that $\bm{m}$, $\bm{n}$, and $\bm{l}=(\hat{\bm{t}}+\hat{\bm{r}})/\sqrt{2}$ form an orthonormal null tetrad). Then, $\phi_{2}\propto E_{{\theta}}+iE_{\phi}$ represents two polarization states of the outgoing electromagnetic waves (in MHD, $E_{{\theta}}$ and $E_{{\phi}}$ correspond to the fast magnetosonic and Alfvén waves, respectively). We have verified that the dominant (quadrupole) component of $\psi_{4}$ observed in our simulation is consistent with the quasi-normal mode computed using qnm code (Stein, 2019). We have also calculated the dominant (dipole) component of $\phi_{2}$, which also approximately matches the corresponding quasi-normal mode frequency (Fig. 4 shows the evolution of dominant spherical harmonics in ${\rm Im}\,\phi_{2}$ and $\psi_{4}$). Note however that $\phi_{2}$ was designed for vacuum electromagnetic fields, and so the oscillation of $\phi_{2}$ may not accurately represent the modulation of MHD outflow. The frequency of compressive modulations observed in Fig. 3 may be directly estimated as $\nu_{\rm mod}=v/\lambda$, where $v\lesssim c$ is the outflow speed and $\lambda\sim 7\,$km is the spatial modulation period. The subsequent evolution of the balding black hole will crucially depend on the resistivity in the current sheet (Bransgrove et al., 2021; Selvi et al., 2024). In the absence of a controlled resistivity in our simulations, we defer the study of the late phase to future work.
3.4 Gamma-ray burst.
@@
Our simulation demonstrates that the magnetospheric destruction during the neutron-star collapse involves strong dissipation and creates a powerful magnetized outflow with a characteristic peak duration $\sim 1$ ms. The hot outflow is launched behind the leading monster shock and has a quasi-periodic tail. The modulated tail is generated by the nascent black hole, as it rings down and quickly sheds most of its magnetic flux initially inherited from the neutron star. The modulation frequency lies naturally in the kilohertz band, suggesting a connection with recently reported kilohertz quasi-periodic oscillations in some GRBs (Chirenti et al., 2023).
@@
Note that the magnetosphere around the neutron star prior to collapse has a minute plasma density. Therefore, the explosion triggered by collapse is practically clean from baryons. The created $e^{\pm}$ plasma in the hot outflow is initially opaque to scattering. Most of the heat density $U$ is contained in trapped blackbody radiation $U\approx aT^{4}$ ($a$ is the radiation constant), as the photons far outnumber the $e^{\pm}$ pairs. The outflow expansion to large radii is not followed by our simulation, however its basic features can be predicted in analogy with the well-known “fireball” model for cosmological GRBs (Paczynski, 1986; Goodman, 1986). The outflow Lorentz factor $\gamma$ will grow and its temperature $T$ will drop due to adiabatic cooling. Eventually, most of $e^{\pm}$ pairs annihilate, and the trapped photons are released, producing a GRB. The burst duration is set by the outflow duration, which lasts only a few ms after the collapse.
@@
The burst spectrum will peak at photon energies $\sim 3k_{B}T\gamma$, where $k_{B}$ is the Boltzmann constant. Simplest GRB models assume adiabatic outflows with no magnetic fields, and then $\gamma T\approx const=T_{0}$ (Paczynski, 1986). In that case, the observed temperature is weakly changed during the outflow expansion and remains close to the initial temperature. Outflows predicted by our simulation are magnetically dominated, and their temperature may be affected by additional dissipation of magnetic energy at large radii. Regulation of the GRB spectral peak in dissipative outflows is discussed in Beloborodov (2013).
@@
The energy of the GRB outflow is set by the pre-collapse energy of the neutron-star magnetosphere ${\cal E}_{m}\sim 10^{47}\mu_{33}^{2}\,$erg. During the collapse the magnetospheric energy is amplified by compression, and then most of it is ejected in the outflow. Assuming that $\sim 0.1$ of this energy is eventually emitted in the GRB, we can roughly estimate the GRB luminosity as $L\sim 0.1{\cal E}_{m}/1\,{\rm ms}\sim 10^{49}\,B_{15}^{2}\,$erg/s, where $B_{15}=B/10^{15}\,\rm G$ is the magnetic field near the neutron star prior to collapse. Note that the outflow (and hence the GRB) is anisotropic, but not strongly collimated, unlike jet-powered GRB models.
@@
During its lifetime prior to collapse, the rotating magnetar produced a magnetic wind outside the dipole magnetosphere, with a termination shock of a large radius behind the merger ejecta (e.g., Metzger et al. 2007; Dessart et al. 2009) 111 Energy deposited by the magnetar wind into the merger ejecta was proposed to power an optical transient much brighter than a usual kilonova (Metzger & Piro, 2014), and non-detection of this effect can constrain the magnetar lifetime (Wang et al., 2023). . The collapse launches a relativistic shock, which will continue to expand into the cold wind. The shock is expected to emit a fast radio burst (FRB) at radii $r\sim 10^{13}$ cm (Beloborodov, 2020). This suggests a mechanism for a delayed FRB from neutron star mergers, emitted together with the delayed GRB. However, the FRB can hardly escape through the surrounding shell of mass $M_{\rm ej}\sim 10^{-2}M_{\odot}$ ejected earlier by the merger (Bhardwaj et al., 2023; Radice et al., 2023). A similar problem is faced by the recently proposed FRB-GRB connection (Rowlinson et al., 2023).
The massive shell is also a threat to the GRB predicted by our simulation. The GRB will not be blocked by $M_{\rm ej}$ if the magnetar collapse occurs with a sufficient delay $\Delta t\gtrsim 10$ hr after the merger, so that the massive ejecta expands to radius $R\sim 3\times 10^{14}\,(\Delta t/10{\rm\,hr})$ cm and its optical depth to gamma-rays $\tau\sim\kappa M_{\rm ej}/4\pi R^{2}$ drops below unity ($\kappa$ is the ejecta opacity to Compton scattering).
4 Acknowledgments
The authors gratefully acknowledge insightful discussions with A. Levinson, M. Lyutikov, A. Philippov, V. Ravi, R. Sari, L. Sironi, J. Stone, S. Teukolsky and C. Thompson. AMB and ERM acknowledge support from NASA’s ATP program under grant 80NSSC24K1229. ERM acknowledges support by the National Science Foundation under grants No. PHY-2309210 and AST-2307394. A.M.B. is supported by NSF AST 2009453, NASA 21-ATP21-0056 and Simons Foundation award No. 446228. B.R. is supported by the Natural Sciences & Engineering Research Council of Canada (NSERC) and by a grant from the Simons Foundation (MP-SCMPS-00001470). Simulations were performed on the NSF Frontera supercomputer under grant AST21006, and on Delta at the National Center for Supercomputing Applications (NCSA) through allocation PHY210074 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.